OpenStudy (anonymous):

Can someone help me? I need someone to tell me if the majority of my answers are right....

jhonyy9 (jhonyy9):

where on what ?

OpenStudy (anonymous):

On 49 questions... I have a feeling that most of them are wrong and this is my final exam... I really don't want to fail it.

jhonyy9 (jhonyy9):

can you rewrite it here ?

OpenStudy (anonymous):

All of my answers and questions? Yes.

OpenStudy (anonymous):

1.Evaluate: -12 + 6 -6 2.Evaluate: -5 - (-2) -3 3.Evaluate: 12 + 10 ÷ 2 - 3 8 4.Simplify the following expression: 8x - 10y + 5z - 5x - 16y + 6z 3x - 6y + 11z 5.Evaluate the following expression using the values given.Find 4x - 6y if x = -2, y = 3 -26

jhonyy9 (jhonyy9):

1,2,3 are right 4 not is correct 5 is right

OpenStudy (anonymous):

I'm pretty sure, 4 is right.

jhonyy9 (jhonyy9):

not is because -10y-16y=-26y

jhonyy9 (jhonyy9):

ok ?

OpenStudy (precal):

3 is wrong 12+10/2-3 12+5-3 17-3 14

OpenStudy (precal):

jhonyy9 is write about 4 it is wrong should be 3x-26y+11z